# Bounds on the number of squares in recurrence sequences

### v2

@@ 3.4. Miscellaneous Lemmas
[…]
@@
$\omega=\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)=e^{i\varphi}$
with $-\pi<\varphi\leq\pi$. Then
$\left|\varphi\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}<0.6.$
(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $y_{\ell}>y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, $\omega$ be as in part (a) and put
@@
with $r_{k},r_{\ell},s_{k},s_{\ell}$ as in Proposition 3.1. Furthermore, suppose that the quantities $\zeta_{4}^{(k)}$ and $\zeta_{4}^{(\ell)}$ defined in (3.20) in the proof of Lemma 3.8 satisfy $\zeta_{4}^{(k)}=\pm\zeta_{4}^{(\ell)}$. Then
$\min_{0\leq j\leq 3}\left|\omega^{1/4}-\zeta_{4}^{j}\frac{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
occurs for either $j=0$ or $j=2$.
@@
(a) We can write $\omega=\left(\left|x_{k}\right|\pm\sqrt{N_{\alpha}}\right)^{2}/\left(x_{k}^{2}+\left|N_{\alpha}\right|\right)$, so with $\omega=e^{i\varphi}$, we have $\left|\tan\left(\varphi\right)\right|=\left|2x_{k}\sqrt{\left|N_{\alpha}\right|}/\left(x_{k}^{2}+N_{\alpha}\right)\right|$. From (1.1) and $N_{\alpha}>-b^{4}d$, we have
@@
From $|\varphi|\leq|\tan(\varphi)|$, we have
$|\varphi|\leq\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{x_{k}^{2}+N_{\alpha}}\right|<\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{0.875x_{k}^{2}}\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}.$
@@
Combining this with the inequality above it yields $|\varphi|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
@@
Using this relationship for $i=\ell$ and using our expressions in Proposition 3.1 for $x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}$ and $f_{k}\sqrt{y_{k}}$, we have
$\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
$-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
@@
(3.29) $|f(x,y)|=$ $\left|\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right.$
$\left.-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
@@
$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
From (3.29), our expression in the statement of this lemma for $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$, the expressions for $y_{k}$ and $y_{\ell}$ in Proposition 3.1, and (1.1) (which implies that $\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|^{2}=x_{k}^{2}-N_{\alpha}=dy_{k}^{2}$), we have
(3.30) $\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$
$=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
@@
(3.31) $\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
@@
From part (a), we have $|\varphi|<0.6$, so $\left|\omega^{1/4}-1\right|<0.15$ and
$\left|\omega^{1/4}-\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|\leq\left|\omega^{1/4}-1\right|+\left|\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-1\right|<0.403.$
@@
$\left|\omega^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|$
$=$ $\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
$\geq\left|\zeta_{4}-\zeta_{4}^{\prime}\right|-\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\sqrt{2}-0.127,$
@@
If $x_{k}$ and $y_{k}$ are both integers, then $\gcd\left(x_{k},y_{k}\right)/\gcd(a,b)=1$ or $2$.
@@
We may assume without loss of generality that $\gcd(a,b)=1$. Otherwise, consider below the sequences $\left(x_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ and $\left(y_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ instead of $\left(x_{k}\right)_{k=0}^{k=\infty}$ and $\left(y_{k}\right)_{k=0}^{k=\infty}$.
@@
$\gcd\left(x_{k},y_{k}\right)\text{ divides }\left\{\begin{array}[]{ll}\gcd\left(x_{0},y_{0}\right)=1&\text{if $k$ is even},\\ \gcd\left(x_{1},y_{1}\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right)=2&\text{if $k$ is odd}.\end{array}\right.$
@@
Put $v_{p}\left(g^{2}\right)=\ell\geq 0$. We will prove that $v_{p}\left(\gcd\left(a^{2},d\right)\right)=\ell$. We also put $v_{p}\left(g_{1}^{2}\right)=\ell_{1}$ and $v_{p}\left(g_{2}\right)=\ell_{2}$. Note that $\ell_{2}=0$ or $1$, since $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$ (from the definition of $g_{2}$, we have $g_{2}|t^{\prime}$ and here $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$). From the definition of $g_{3}$, we have $g_{3}|4$, so $v_{p}\left(g_{3}\right)=0$ for odd primes, $p$. Hence $\ell=\ell_{1}+\ell_{2}$.
@@
In the same way, we find that $p^{\ell_{1}+2\ell_{2}}|x_{k}^{2}$.
By Lemma 3.10, since $p$ is odd, $p\nmid\gcd\left(x_{k},y_{k}\right)$. Hence, from $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $p^{\ell}|d$. Combining this with $p^{\ell}|N_{\alpha}$ and $N_{\alpha}=a^{2}-db^{4}$, we find that $p^{\ell}|a^{2}$. So $p^{\ell}|\gcd\left(a^{2},d\right)$.
@@
[…]
@@
Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, $k=-1$, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$.
@@
For $d\geq 155$, we find that $1.130.091\sqrt{d}>$ 1.13 holds.
@@
where we use $|g|\mathcal{N}_{d^{\prime},4}\leq 8\sqrt{\left|N_{\alpha}\right|}$ from (3.36) and Lemma 3.5(c) with $u\geq 1$ to establish the first inequality. For $d\geq 109,000$, we find that $2170.66\sqrt{d}>$ 217 holds.
@@
For each pair, we check (3.37) and (3.38) by computing their left-hand sides for all $k\neq 0$ such that $20.1832\cdot 2^{1+\min\left(v_{2}(b)/2,2\right)}\sqrt{d}y_{k}/b<$ 2 and $30021.12y_{k}/2^{1+\min\left(v_{2}(b)/2,2\right)}<$ 300. No counterexamples to (3.37) and (3.38) were found. A PARI/GP program took 112 seconds to run on a Windows 10 laptop with an Intel i7-9750H processor and 16GB of RAM.
The smallest value of the left-hand side of (3.37) found for $d\geq 105$ was $1.139\ldots$, for $a=11$, $d=140$, $N_{\alpha}=-19$, $k=-1$, $x_{k}=-59$ and $y_{k}=5$. The smallest value of the left-hand side of (3.38) found for $d\geq 105$ was $217.3\ldots$, for $a=10$, $d=140$, $N_{\alpha}=-40$, $k=-1$, $x_{k}=-130$ and $y_{k}=11$.
@@
If $b=1$ and $-N_{\alpha}$ is a square, then there is at most one integer square among all distinct elements of $\left(y_{k}\right)_{k=-\infty}^{\infty}$ which satisfies
@@
Conjecture 1.2 would immediately follow when $-N_{\alpha}$ is a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$. This is due to Lemma 3.5.

### v3

@@ 3.4. Miscellaneous Lemmas
Let the $y_{k}$’s be defined as in (1.1) with the notation and assumptions there. Suppose that $N_{\alpha}<0$.
@@
$\omega_{k}=\left(x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)=e^{i\varphi_{k}}$
with $\pi<\varphi_{k}\leq\pi$. Then
$\left|\varphi_{k}\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}<0.6.$
(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $k,\ell\neq 0$, $y_{\ell}>y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, $\omega_{k}$ be as in part (a) and put
@@
with $_{k}, rr_{\ell}$, $s_{k}$ and $s_{\ell}$ as in Proposition 3.1. Furthermore, suppose that the quantities $\zeta_{4}^{(k)}$ and $\zeta_{4}^{(\ell)}$ defined in (3.20) in the proof of Lemma 3.8 satisfy $\zeta_{4}^{(k)}=\pm\zeta_{4}^{(\ell)}$. Then
$\min_{0\leq j\leq 3}\left|\omega_{k}^{1/4}-\zeta_{4}^{j}\frac{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
occurs for either $j=0$ or $j=2$, where $\zeta_{4}$ is a primitive $4$-th root of unity.
@@
(a) We can write $\omega_{k}=\left(\left|x_{k}\right|\pm\sqrt{N_{\alpha}}\right)^{2}/\left(x_{k}^{2}+\left|N_{\alpha}\right|\right)$, so with $\omega_{k}=e^{i\varphi_{k}}$, we have $\left|\tan\left(\varphi_{k}\right)\right|=\left|2x_{k}\sqrt{\left|N_{\alpha}\right|}/\left(x_{k}^{2}+N_{\alpha}\right)\right|$. From (1.1) and $N_{\alpha}>-b^{4}d$, we have
@@
From $\left|\varphi_{k}\right|\leq\left|\tan\left(\varphi_{k}\right)\right|$, we have
$\left|\varphi_{k}\right|\leq\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{x_{k}^{2}+N_{\alpha}}\right|<\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{0.875x_{k}^{2}}\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}.$
@@
Combining this with the inequality above it yields $\left|\varphi_{k}\right|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
@@
Using this relationship for $i=\ell$ and using our expressions in Proposition 3.1 for $x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}$ and $f_{k}\sqrt{y_{k}}$, we have
$\left(x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
$-\left(x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
@@
(3.29) $|f(x,y)|=$ $\left|\left(x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right$.
$\left.-\left(x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
@@
$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega_{k}^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
From (3.29), our expression in the statement of this lemma for $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$, the expressions for $y_{k}$ and $y_{\ell}$ in Proposition 3.1, and (1.1) (which implies that $\left|x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right|^{2}=x_{k}^{2}-N_{\alpha}=dy_{k}^{2}$), we have
(3.30) $\left|\omega_{k}-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$
$=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}\left|x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
@@
(3.31) $\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.1263.$
@@
From part (a), we have $\left|\varphi_{k}\right|<0.6$, so $\left|\omega_{k}^{1/4}-1\right|<0.15$ and
$\left|\omega_{k}^{1/4}-\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|\leq\left|\omega_{k}^{1/4}-1\right|+\left|\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-1\right|<0.403.$
@@
$\left|\omega_{k}^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|$
$=$ $\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
$\geq\left|\zeta_{4}-\zeta_{4}^{\prime}\right|-\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\sqrt{2}-0.127,$
@@
If $x_{k}$ and $y_{k}$ are both integers and $\gcd(a,b)$ is odd, then $\gcd\left(x_{k},y_{k}\right)/\gcd(a,b)=1$ or $2$.
The additional hypothesis that $\gcd(a,b)$ is odd was omitted from the published version of this lemma. If $\gcd(a,b)$ is even, then $\gcd\left(x_{k},y_{k}\right)/\gcd(a,b)=1/2$ also occurs.
@@
We may assume without loss of generality that $\gcd(a,b)=1$. Otherwise, consider below the sequences $\left(x_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ and $\left(y_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ instead of $\left(x_{k}\right)_{k=0}^{k=\infty}$ and $\left(y_{k}\right)_{k=0}^{k=\infty}$. Since $x_{0}=a$, $x_{1}=\left(at+bdu\right)/2$, $x_{-1}=\pm\left(at-dbu\right)/2$, $y_{0}=b$, $y_{1}=\left(au+bt\right)/2$, $y_{-1}=\pm\left(au-bt\right)/2$, the recurrence relation satisfied by both sequences has integer coefficients and $\gcd(a,b)$ is odd, we see that $x_{k}/\gcd(a,b)$ and $y_{k}/\gcd(a,b)$ are integers if and only if $x_{k}$ and $y_{k}$ are integers.
@@
[…]
@@
Put $v_{p}\left(g^{2}\right)=v_{p}\left(g_{1}^{2}g_{2}/g_{3}\right)=\ell$. Since $g_{3}$ is a power of $2$, it follows that $v_{p}\left(g_{3}\right)=0$ and hence $\ell\geq 0$. We will prove that $v_{p}\left(\gcd\left(a^{2},d\right)\right)=\ell$. We also put $v_{p}\left(g_{1}^{2}\right)=\ell_{1}$ and $v_{p}\left(g_{2}\right)=\ell_{2}$. Note that $\ell_{2}=0$ or $1$, since $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$ (from the definition of $g_{2}$, we have $g_{2}|t^{\prime}$ and here $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$). Since $v_{p}\left(g_{3}\right)=0$, we have $\ell=\ell_{1}+\ell_{2}$.
@@
In the same way, we find that $p^{\ell_{1}+2\ell_{2}}|x_{k}^{2}$. So $p^{\ell}|x_{k}^{2}$ too.
By Lemma 3.10 and $\gcd(a,b)=1$, since $p$ is odd, $p\nmid\gcd\left(x_{k},y_{k}\right)$. Hence, from $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $p^{\ell}|d$. Combining this with $p^{\ell}|N_{\alpha}$ and $N_{\alpha}=a^{2}-db^{4}$, we find that $p^{\ell}|a^{2}$. So $p^{\ell}|\gcd\left(a^{2},d\right)$.
@@
Suppose that $a$, $b$, $d$, $g$, $t^{\prime}$, $u_{1}$, $u_{2}$, $x_{k}$ and $y_{k}$ are as in Lemma 3.11, $d^{\prime}=u_{2}^{2}t^{\prime}/g^{2}$, as defined in (2.10), $\mathcal{N}_{d^{\prime},4}$ is as defined in (2.1) and put $b^{\prime}=N_{\alpha}/\gcd\left(a^{2},db^{4}\right)$. Then
(3.36) $|g|\mathcal{N}_{d^{\prime},4}=\sqrt{\gcd\left(a^{2},db^{4}\right)}\,2^{1+\min\left(2,v_{2}(b^{\prime})/2\right)}\geq 2^{1+\min\left(2,v_{2}\left(N_{\alpha}\right)/2\right)}.$
@@
Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, k=1k=-1, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$.
@@
For $d\geq 155$, we find that $0.091\sqrt{d}>1.13$ holds.
@@
where we use $|g|\mathcal{N}_{d^{\prime},4}\leq 8\sqrt{\left|N_{\alpha}\right|}$ from (3.36) and Lemma 3.5(c) with $u\geq 1$ to establish the first inequality. For $d\geq 109,000$, we find that $0.66\sqrt{d}>217$ holds.
@@
For each pair, we check (3.37) and (3.38) by computing their left-hand sides for all $k\neq 0$ such that $0.1832\cdot 2^{1+\min\left(v_{2}(b)/2,2\right)}\sqrt{d}y_{k}/b<2$ and $21.12y_{k}/2^{1+\min\left(v_{2}(b)/2,2\right)}<300$. No counterexamples to (3.37) and (3.38) were found. A PARI/GP program took 112 seconds to run on a Windows 10 laptop with an Intel i7-9750H processor and 16 GB of RAM.
The smallest value of the left-hand side of (3.37) found for $d\geq 105$ was $1.139\ldots$, for $a=11$, $d=140$, $N_{\alpha}=-19$, k=1k=-1, $x_{k}=-59$ and $y_{k}=5$. The smallest value of the left-hand side of (3.38) found for $d\geq 105$ was $217.3\ldots$, for $a=10$, $d=140$, $N_{\alpha}=-40$, k=1k=-1, $x_{k}=-130$ and $y_{k}=11$.
@@
If $b=1$ and $N_{\alpha}$ is a square, then there is at most one integer square among all distinct elements of $\left(y_{k}\right)_{k=-\infty}^{\infty}$ which satisfies
@@
Conjecture 1.2 would immediately follow when $N_{\alpha}$ is a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$. This is due to Lemma 3.5.